S_{2} - S_{0} FLUORESCENCE, MOLECULAR STRUCTURE, AB INITIO CALCULATIONS, AND QUANTUM BEAT SPECTROSCOPY OF SILYLIDENE (H_{2}C=Si)

\underline{WARREN W. HARPER}, KEVIN W. WADDELL, RANDALL K. HILLIARD, ROGER S. GREV, AND DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington KY 40506-0055.

Strong S_{2} - S_{0} fluorescence has been observed in LIF spectra of jet-cooled silylidene, H_{2}C=Si, produced by fragmentation of tetramethylsilane in a pulsed discharge jet. The experimental ground and excited state molecular structures have been obtained for the first time from rotational analyses of the 0_{0}^{0} bands of H_{2}C=Si and D_{2}C=Si. These bands exhibit pronounced rotational level intensity anomalies indicative of nonradiative decay processes. Quantum beats are observed in the fluorescence decay of almost every rotational level in v′ = 0 and in most of the higher vibrational levels. Fourier transforms of the beat patterns show line broadening and splittings at small magnetic fields of 10 - 70 Gauss. High quality ab initio studies have been done to locate the interacting electronic states and predict their vibrational frequencies and molecular structures.